Given that u(t)=2tē-1, find the expression for u(3t)-2u(t)
Answer by oscargut(891)

u(t)=2tē-1, find the expression for u(3t)-2u(t)
u(3t)=2(3t)^2-1=18t^2-1
u(2t)=2(2t)^2-1=8t^2-1
then
u(3t)-2u(t)=18t^2-1-2(8t^2-1)=2t^2+1
Answer:u(3t)-2u(t)=2t^2+1